You have been asked to develop a capitation rate for a primary care group based on the following projections: Service Annual Frequency/1,000 Cost per Service Inpatient Visits 100 $7,000.00 Office Visits 3,000 $45.00 Lab/X-ray 500 $25.00 What per-member per-month (PMPM) rate would be required to break even, ignoring any copayments?

A. $71.7980

B. $70.6250

C. $58.3333

. Not enough data to solve
Answer #1

To calculate PMPM, first of all we shall determine the expected encounter per patient per year (annual per patient utilization rate) as following

Inpatient visit utilization/member= 100/1000 = 0.1 inpatient visits per member per year

Office visit utilization/member= 3,000/1,000= 3.0 outpatient visits per member per year

Lab/X-ray utilization/member= 500/1,000= 0.5 labs/x-rays per member per year

Now we shall calculate the cost per member per year, than divide it by 12 to get PMPM

Inpatient visits

= 0.1 x $7,000

= $700.00

= $ 700/12

= $58.3333 Approximately

Office visits

= 3.0 x $ 45

= $135.00

= $ 135/12

= $11.25

Lab/x-rays

= 0.5 x $ 25

= $ 12.50

= $ 12.5/12

= $ 1.0417 Approximately

So the per-member per-month (PMPM) rate that would be required to break even will be

= $ 58.3333 + $ 11.25 + $ 1.0417

= $ 70.6250 Approximately

So the correct answer is option B i.e. $ 70.6250

Feel free to ask in case of any query relating to this question